Can someone help me on this?
I need to solve: 7x - 5 = 6x
Thanks,
Lyn

Answer by wgomero(28)

Resolution:
7x - 5 = 6x; move 6x to the left side by substracting (6x) both sides
7x - 6x - 5 = 6x - 6x
x -5 = 0
move 5 to the right side by adding (5) both sides
x - 5 + 5 = 5
therefore x = 5